The total price of 4 pens and 2 copies is rs 160 and that pen
Of 5 pen and 6 copies is rs 270 . Find the price of a single pen and single copy

Answer by ikleyn(52788)

.

    4p + 2c = 160     (1)

    5p + 6c = 270     (2)


Multiply equation (1) by 3 (both sides).  Keep equation (2) as is.


    12p + 6c = 480    (1')

     5p + 6c = 270    (2')


From equation (1') subtract equation (1').  Doing this way, you eliminate the terms "6c", which cancel each other.


    12p - 5p = 480 - 270

        7p   = 210

         p   = 210/7 = 30.


To find "c", substitute the found value p= 30 into equation (1)


    4*30 + 2c = 160

    120  + 2c = 160

          2c  = 160 - 120 = 40

           c  = 40/2 = 20.


ANSWER.  A pen price is rs 30;  a copy price is  rs 20 for each.

Solved.

The method I used in my solution is called "the Elimination method".

You may check my solution on your own by substituting the found values into the original equations.